OpenStudy (erinecookie):

Write the equation of a line with a slope of –1 and a y-intercept of –6.

OpenStudy (erinecookie):

y = –x – 6 y = –6x – 1 –x + y = –6 –x – y = –6 are the answer choices

OpenStudy (niksbhalla14mar):

Slope is -1 and y intercept means when x=0, then y =-6 , second co-ordinate would be the x intercept where y=0 , then x would be some value. Co-ordinates are (0,-6) and (x,0). Slope = (y2-y1)/(x1-x2) = (0+6)/(0-x) =-1 So. x = 6 when y =0 So option c is correct -x+y=-6 as it holds good for both (0,-6) and (6,0)

OpenStudy (erinecookie):

Thank you, that makes sense

OpenStudy (erinecookie):

can you help me with another? @niksbhalla14mar